A volume of 450 cubic centimeter is measured at a pressure of 740 mm Hg and a temperature of 25.4 oC.


What is the volume in liter? Pressure in pound per square inch and temperature in degree Kelvin?


1
Expert's answer
2021-07-07T05:03:08-0400

Given Volume = 450 cc


Pressure = 740 mm Hg


Temperature = 25 . 4"\degree" C



Volume in liters = "\dfrac{volume \ in \ cubic \ centimetres}{1000}" = "\dfrac{450}{1000}"



Volume in liters = 0. 45 ltr




Pressure in pound per square inch = "\dfrac{Pressure \ in \ mm \ Hg}{0.019}" = "\dfrac{740}{0.019}"



Pressure in pound per square inch = 38947.36 psi




Temperature in kelvin = Temperature in "\degree"C + 273 = 25.4 + 273


Temperature in kelvin = 298.4 kelvin